OpenStudy (anonymous):

what is 1/4 of 3200 and 1400???

OpenStudy (anonymous):

3200 * 0.25 1400 * 0.25

OpenStudy (anonymous):

thank you very much:D its for a geometry class i wasent sure wether i was right or wrong:)

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

if i had to turn this into a swimming pool,would this turn out to be 350x800 or should i switch it?

OpenStudy (anonymous):

the order of your numbers shouldn't matter.

OpenStudy (anonymous):

ah ok im just making sure cause i really really want to pass this project:) thanks again:)

OpenStudy (anonymous):

Give me some more details and maybe I can send you off in the right direction

OpenStudy (anonymous):

well the first part was Mrs. Noether wants the pool to hold between 1400 ft3 and 3200 ft3 of water and be from 6-10 feet deep. then you have to 1.The dimensions ,2. amount of pool liner material that will be needed to completely line the inside of the pool (sides and bottom) and then 3.3.The amount of water that will be needed to fill the pool if the water needs to be 6 inches below the top of the pool

OpenStudy (anonymous):

This is a volume related question. Do you know the volume for a rectangle?

OpenStudy (anonymous):

Seems like you get to pick the dimensions of the pool more or less. You're told the height of it has to be between 6-10 feet deep. However the length and width could be anything as long as the volume is still between 1400ft3 and 3200ft3

OpenStudy (anonymous):

yeah i did smallest:if held 1400ft3,10ft deep since v=lwh low width=1400/10=140 largest:if held 3200ft3,was only 6ft deep then 3200/6= around 533 or 53.3 when you move over the decimal

OpenStudy (anonymous):

but now im having trouble doing the dog house which has to be 1/4 the size of the bigger pool

OpenStudy (anonymous):

v=lwh 1400 = lw(10) lw = 140 (this is area) Your l could = 10 and w = 14. From here you need to find the areas of all the sides and then find the amount of material. Remember that a pool has 4 sides and a bottom. Try sketching it out with the dimensions..it'll help to visualize it.

OpenStudy (anonymous):

yeah i drew it out on the paper to make sure i did it right cause it is easier to visualize then to do it all in your mind like einstien:)